### v2

# On the free metabelian Novikov and metabelian Lie-admissible algebras

Abstract.
In this paper, we consider Lie-admissible algebras, which are free Novikov and free Lie-admissible algebras with an additional metabelian identity. We construct a linear basis for both free metabelian Novikov and free metabelian Lie-admissible algebras. Additionally, we describe a space of symmetric polynomials for both the free metabelian Novikov algebra and the free metabelian Lie-admissible algebra.
Key words and phrases:
Novikov algebra, Lie-admissible algebra, metabelian identity, free algebra, polynomial identities
2010 Mathematics Subject Classification:
17A30, 17A50, 17D25
1. Introduction
In recent years, algebras with metabelian identity have become popular objects in ring theory. Metabelian identity also can be stated as the solvability of index $2$. Various types of classical algebras with metabelian identity, such as Lie, Leibniz, Malcev, Jordan, etc. are considered. The variety of metabelian Lie algebras has attracted significant attention, see [17, 22]. A basis of the free metabelian Lie algebra was constructed in [1]. In an analogical way, a basis of the free metabelian Leibniz algebra was constructed in [8]. Symmetric polynomials in the free metabelian Lie algebras were considered in [9]. The generators of symmetric polynomials in free metabelian Leibniz algebras were found in [21]. Other examples of Lie-admissible algebras are assosymmetric algebras [6]. A basis of the free metabelian Malcev algebra is constructed in [20]. For this reason, we add metabelian identity to a well-known class of algebras which are Novikov and Lie-admissible.
Novikov algebras were introduced in a study of Hamiltonian operators concerning the integrability of certain partial differential equations [12]. Later, they played a significant role in a study of Poisson brackets of hydrodynamic type [2].
It is well-known that, given a commutative algebra $A$ with a derivation $D$, the space $A$ under the product $x_{1}\circ x_{2}=D(x_{1})x_{2}$ forms a (right) Novikov algebra. Moreover, every Novikov algebra can be embedded into an appropriate commutative algebra with derivation $D$ [4]. Using rooted trees, the monomial basis of the free Novikov algebra in terms of $\circ$ was constructed in [11]. In terms of Young diagrams, the basis was developed in [10]. By utilizing commutative algebra with derivation $D$ and a well-defined order, an alternative monomial basis of the free Novikov algebra is presented in [15]. The issues of solvability and nilpotency of Novikov algebras were addressed in [24]. In section $2$, we construct a basis of the free solvable Novikov algebra with an index of $2$. In section $3$, we explicitly describe the symmetric polynomials for the multilinear part of the free metabelian Novikov algebra.
Let’s shift our focus to metabelian Lie-admissible algebras. In the realm of Lie-admissible algebra theory, additional conditions like flexibility or power-associativity are crucial, leading to numerous noteworthy outcomes in this context, see [3, 19]. Another important direction of the research on Lie-admissible algebras concerns the property that their associator satisfies relations defined by a natural action of the symmetric group of degree 3 [13, 14]. A basis of the free Lie-admissible algebra and the Gröbner-Shirshov base theory for Lie-admissible algebras is given in [5]. In addition, there is given an analogue of PBW-theorem for the pair of Lie and Lie-admissible algebras. A basis of the free Lie-admissible algebra in terms of commutator and anti-commutator is given in [18]. In section $4$, we construct the basis of a free metabelian Lie-admissible algebra in terms of commutator and anti-commutator. In Section $5$, we explicitly describe the symmetric polynomials for the multilinear part of the free metabelian Lie-admissible algebra.
We consider all algebras over a field $\mathbb{K}$ of characteristic 0.
2. Free metabelian Novikov algebra
An algebra is called metabelian (right) Novikov if it satisfies the following identities:
$a(bc)=b(ac),$ (1)
$(ab)c-a(bc)=(ac)b-a(cb),$ (2)
$(ab)(cd)=0$ (3)
Let $X=\{x_{1},x_{2}\ldots\}$ be a countable set of generators. We denote by $\mathrm{Nov}\langle X\rangle$ and $\mathrm{M}\mathrm{Nov}\langle X\rangle$ the free Novikov and free metabelian Novikov algebra, respectively.
Let us denote by $\mathcal{N}_{n}$ the set of monomials of degree $n$ of the following form:
$\mathcal{N}_{n}=\{(\cdots((x_{i_{1}}x_{i_{2}})x_{i_{3}})\cdots)x_{i_{n}}\;|\;i_{2}\leq i_{3}\leq\ldots\leq i_{n}\},$
where $n\geq 5$. We set
$\mathcal{N}_{1}=\{x_{i}\},\mathcal{N}_{2}=\{x_{i_{1}}x_{i_{2}}\}.$
For degree $3$, $\mathcal{N}_{3}$ is defined as follows:
$\mathcal{N}_{3}=\{(x_{i_{1}}x_{i_{2}})x_{i_{3}},x_{i_{3}}(x_{i_{2}}x_{i_{1}})\;|\;i_{2}\leq i_{3}\}.$
For degree $4$, we define the set $\mathcal{N}_{4}$ as follows:
$\mathcal{N}_{4}=\{((x_{i_{1}}x_{i_{2}})x_{i_{3}})x_{i_{4}},x_{j_{1}}(x_{j_{2}}(x_{j_{3}}x_{j_{4}}))\;|\;i_{2}\leq i_{3}\leq i_{4},\;j_{1}\leq j_{2}\leq j_{3}\leq j_{4}\}.$
2.1 Lemma.
The algebra $\mathrm{M}\mathrm{Nov}\langle X\rangle$ satisfies the following identities:
$((x_{1}(x_{2}x_{3}))x_{4})x_{5}=(x_{1}((x_{2}x_{3})x_{4}))x_{5}=(x_{1}(x_{2}(x_{3}x_{4})))x_{5}=x_{1}(((x_{2}x_{3})x_{4})x_{5})=\\ x_{1}((x_{2}(x_{3}x_{4}))x_{5})=x_{1}(x_{2}((x_{3}x_{4})x_{5}))=x_{1}(x_{2}(x_{3}(x_{4}x_{5})))=0.$ (4)
Proof.
Using (1), (2) and (3), one can obtain
$0=((x_{1}x_{2})x_{3})(x_{4}x_{5})=x_{4}(((x_{1}x_{2})x_{3})x_{5})=(x_{1}(x_{2}x_{5}))(x_{4}x_{3})=x_{4}((x_{1}(x_{2}x_{5}))x_{3})=\\ x_{4}(x_{1}((x_{2}x_{5})x_{3}))+x_{4}((x_{1}x_{3})(x_{2}x_{5}))-x_{4}(x_{1}(x_{3}(x_{2}x_{5})))=\\ x_{4}((x_{2}x_{5})(x_{1}x_{3}))-x_{4}(x_{1}(x_{3}(x_{2}x_{5})))=-x_{4}(x_{1}(x_{3}(x_{2}x_{5}))),$
which gives
$x_{1}(((x_{2}x_{3})x_{4})x_{5})=x_{1}((x_{2}(x_{3}x_{4}))x_{5})=x_{1}(x_{2}(x_{3}(x_{4}x_{5})))=x_{1}(x_{2}((x_{3}x_{4})x_{5}))=0.$
By (1), (2) and (3), we have
$((x_{1}(x_{2}x_{3}))x_{4})x_{5}=(x_{1}((x_{2}x_{3})x_{4}))x_{5}+((x_{1}x_{4})(x_{2}x_{3}))x_{5}-(x_{1}(x_{4}(x_{2}x_{3})))x_{5}=\\ -x_{1}((x_{4}(x_{2}x_{3}))x_{5})-(x_{1}x_{5})(x_{4}(x_{2}x_{3}))+x_{1}(x_{5}(x_{4}(x_{2}x_{3})))=x_{1}(x_{5}(x_{4}(x_{2}x_{3})))=\\ -(x_{4}(x_{2}x_{3}))(x_{1}x_{5})+x_{1}(x_{5}(x_{4}(x_{2}x_{3})))=x_{1}(x_{5}(x_{4}(x_{2}x_{3})))=0,$
which gives
$((x_{1}(x_{2}x_{3}))x_{4})x_{5}=(x_{1}((x_{2}x_{3})x_{4}))x_{5}=(x_{1}(x_{2}(x_{3}x_{4})))x_{5}=0.$
2.2 Theorem.
The set $\bigcup\mathcal{N}_{i}$ is a linear basis of the algebra $\mathrm{M}\mathrm{Nov}\langle X\rangle$.
Proof.
The statement for degrees less than 5 can be verified by direct calculations. For $n\geq 5$, firstly, we show that every monomial of $\mathrm{M}\mathrm{Nov}\langle X\rangle$ can be written as a sum of monomials from the set $\bigcup_{i\geq 5}\mathcal{N}_{i}$. By Lemma (2.1), we obtain that every monomial except $(\cdots((x_{i_{1}}x_{i_{2}})x_{i_{3}})\cdots)x_{i_{n}}$ is equal to 00. By (2) and Lemma (2.1), we have
$(\cdots((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{m-1}})x_{i_{m}})\cdots)x_{n}=(\cdots((\cdots(x_{i_{1}}x_{i_{2}})\cdots)(x_{i_{m-1}}x_{i_{m}}))\cdots)x_{n}+\\ (\cdots(((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{m}})x_{i_{m-1}})\cdots)x_{n}-(\cdots((\cdots(x_{i_{1}}x_{i_{2}})\cdots)(x_{i_{m}}x_{i_{m-1}}))\cdots)x_{n}=\\ (\cdots(((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{m}})x_{i_{m-1}})\cdots)x_{n},$
i.e., the generators $x_{i_{2}}$, $x_{i_{3}},\ldots$, $x_{i_{n}}$ are rearrangeable. From these equations, we get that every monomial of $\mathrm{M}\mathrm{Nov}\langle X\rangle$ which has a degree greater than 4 can be written as a sum of $\bigcup_{i\geq 5}\mathcal{N}_{i}$.
Now, we consider an algebra $A\langle X\rangle$ with a basis monomials $\bigcup\mathcal{N}_{i}$ and multiplication *. Let us define a multiplication on monomials $\bigcup\mathcal{N}_{i}$ in $A\langle X\rangle$ as follows:
$\begin{gathered}\begin{cases}X_{1}*X_{2}=0\;\;\text{if $X_{1},X_{2}\in\mathcal{N}_{i}$ and $\mathrm{deg}(X_{1}),\mathrm{deg}(X_{2})>1$};\\ x_{j}*((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{n}})=0;\\ ((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{n}})*x_{j}=((\cdots(((\cdots(x_{i_{1}}x_{i_{2}})\cdots)x_{i_{k}})x_{j})x_{i_{k+1}})\cdots)x_{i_{n}},\\ \end{cases}\end{gathered}$
where $n>4$ and $i_{2}\leq\ldots\leq i_{k}\leq j\leq i_{k+1}\leq\ldots\leq i_{n}$. Up to degree 4, we define multiplication in $A\langle X\rangle$ that is consistent with identities (1), (2) and (3). By straightforward calculation, we can check that an algebra $A\langle X\rangle$ satisfies to (1), (2) and (3) identities. It remains to note that $A\langle X\rangle\cong\mathrm{M}\mathrm{Nov}\langle X\rangle$. ∎
3. Symmetric polynomials of the free metabelian Novikov algebra
Let $p(x_{1},x_{2},\ldots,x_{n})$ be a polynomial of the free metabelian Novikov algebra generated by a finite set $X=\{x_{1},x_{2},\ldots,x_{n}\}$. The polynomial $p(x_{1},x_{2},\ldots,x_{n})$ is called symmetric if it satisfies the following condition:
$\sigma p(x_{1},x_{2},\ldots,x_{n})=p(x_{\sigma(1)},x_{\sigma(2)},\ldots,x_{\sigma(n)})=p(x_{1},x_{2},\ldots,x_{n}),$
where $\sigma\in S_{n}$. Let us define a set of polynomials $\mathcal{P}$ in $\mathrm{M}\mathrm{Nov}\langle X\rangle$ as follows:
$p_{1}=\sum_{i}x_{i},\;p_{2}=\sum_{i\neq j}x_{i}x_{j},$
$p_{3,1}=\sum_{i=1}^{n}\sum_{j_{1}<j_{2}}x_{j_{2}}(x_{j_{1}}x_{i}),\;p_{3,2}=\sum_{i=1}^{n}\sum_{j_{1}<j_{2}}(x_{i}x_{j_{1}})x_{j_{2}},$
$p_{4,1}=\sum_{j_{1}<j_{2}<j_{3}<j_{4}}x_{j_{1}}(x_{j_{2}}(x_{j_{3}}x_{j_{4}})),\;p_{4,2}=\sum_{i=1}^{n}\sum_{j_{1}<j_{2}<j_{3}}((x_{i}x_{j_{1}})x_{j_{2}})x_{j_{3}},$
and
$p_{n}=\sum_{i=1}^{n}\sum_{j_{1}<j_{2}<\ldots<j_{n-1}}(\cdots((x_{i}x_{j_{1}})x_{j_{2}})\cdots)x_{j_{n-1}},$
where $n\geq 5$. The multilinear part of the free metabelian Novikov algebra is a space consisting of all elements containing each $x_{i}$ exactly once.
3.1 Example.
For multilinear part of $\mathrm{M}\mathrm{Nov}\langle X\rangle$, we obtain
$p_{3,1}=x_{2}(x_{1}x_{3})+x_{3}(x_{2}x_{1})+x_{3}(x_{1}x_{2}),\;p_{3,2}=(x_{1}x_{2})x_{3}+(x_{2}x_{1})x_{3}+(x_{3}x_{1})x_{2},$
$p_{4,1}=x_{1}(x_{2}(x_{3}x_{4})),\;p_{4,2}=((x_{1}x_{2})x_{3})x_{4}+((x_{2}x_{1})x_{3})x_{4}+((x_{3}x_{1})x_{2})x_{4}+((x_{4}x_{1})x_{2})x_{3},$
$p_{n}=(\cdots(((x_{1}x_{2})x_{3})x_{4})\cdots)x_{n}+(\cdots(((x_{2}x_{1})x_{3})x_{4})\cdots)x_{n}+\\ (\cdots(((x_{3}x_{1})x_{2})x_{4})\cdots)x_{n}+\ldots+(\cdots(((x_{n}x_{1})x_{2})x_{3})\cdots)x_{n-1},$
where $n\geq 5$.
3.2 Theorem.
For the multilinear part of the free metabelian Novikov algebra, the symmetric polynomials have the form $\mathcal{P}$.
Proof.
For $n=1,2$, the result is obvious. For $n\geq 3$, we use the fact that every Novikov algebra is embeddable into commutative algebra with derivation and spanning elements of Novikov algebra in commutative algebra with derivation are monomials of weight $-1$ [4, 16]. The weight function $\mathrm{wt}(u)\in\mathbb{Z}$ is defined on monomials of commutative algebra with derivation $D$ by induction as follows,
$\mathrm{wt}(x)=-1,\quad x\in X;$
$\mathrm{wt}(d(u))=\mathrm{wt}(u)+1;\quad\mathrm{wt}(uv)=\mathrm{wt}(u)+\mathrm{wt}(v).$
For simplicity, we denote $D(x)$ and $D^{n}(x)$ by $x^{\prime}$ and $(D^{n-1}(x))^{\prime}$, respectively. For the multilinear part to describe the space of symmetric polynomials of $\mathrm{Nov}\langle X\rangle$ in degree $3$, we need to find the space of the symmetric polynomials of the differential commutative algebra of weight $-1$ in degree $3$. This is the monomials of the form
$x_{2}^{\prime}x_{1}^{\prime}x_{3}+x_{3}^{\prime}x_{2}^{\prime}x_{1}+x_{3}^{\prime}x_{1}^{\prime}x_{2}\;\textrm{and}\;x_{1}^{\prime\prime}x_{2}x_{3}+x_{2}^{\prime\prime}x_{1}x_{3}+x_{3}^{\prime\prime}x_{1}x_{2}.$
Rewriting the first polynomial by the operation of Novikov algebra, we obtain $p_{3,1}$. Rewriting the second polynomial, we obtain
$(x_{1}x_{2})x_{3}-x_{1}(x_{2}x_{3})+(x_{2}x_{1})x_{3}-x_{2}(x_{1}x_{3})+(x_{3}x_{1})x_{2}-x_{3}(x_{1}x_{2}).$
Adding to the last expression $p_{3,1}$, we obtain $p_{3,2}$.
For degree $4$, we have $\mathrm{M}\mathrm{Nov}\langle X\rangle/\{(ab)(cd)\}\cong\mathrm{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$, where $\mathrm{Com}\langle X\rangle^{(D)}_{-1}$ is a space of differential commutative algebra of weight $-1$. Hence, in $\mathrm{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ we rewrite all monomials of the form $x_{i_{1}}^{\prime\prime}x_{i_{2}}^{\prime}x_{i_{3}}x_{i_{4}}$ to $-x_{i_{1}}^{\prime}x_{i_{2}}^{\prime}x_{i_{3}}^{\prime}x_{i_{4}}$. It remains to note that
$0=a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d-(a^{\prime\prime}b^{\prime}dc+a^{\prime}b^{\prime}d^{\prime}c)=a^{\prime}b^{\prime}c^{\prime}d-a^{\prime}b^{\prime}d^{\prime}c$
which gives that we rewrite monomial $a^{\prime}b^{\prime}d^{\prime}c$ to $a^{\prime}b^{\prime}c^{\prime}d$. Finally, the remained monomials of weight $-1$ are $x_{i_{1}}x_{i_{2}}x_{i_{3}}x_{i_{4}}^{\prime\prime\prime}$ and $x_{j_{1}}x_{j_{2}}^{\prime}x_{j_{3}}^{\prime}x_{j_{4}}^{\prime}$, where $i_{1}\leq i_{2}\leq i_{3}$ and $j_{1}\leq j_{2}\leq j_{3}\leq j_{4}$. By Theorem 2.2, these monomials are linearly independent, and for multilinear part symmetric polynomials of $\mathrm{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ are
$x_{1}^{\prime}x_{2}^{\prime}x_{3}^{\prime}x_{4}\;\textrm{and}\;x_{1}^{\prime\prime\prime}x_{2}x_{3}x_{4}+x_{2}^{\prime\prime\prime}x_{1}x_{3}x_{4}+x_{3}^{\prime\prime\prime}x_{1}x_{2}x_{4}+x_{4}^{\prime\prime\prime}x_{1}x_{2}x_{3},$
which correspond to $p_{4,1}$ and $p_{4,2}$, analogically as in degree $3$.
By Theorem 2.2, starting from degree $5$, all monomials of $\mathrm{Com}\langle X\rangle^{(D)}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ except $x_{i}^{(n-1)}x_{j_{n-1}}\ldots x_{j_{1}}$ are equal to 00, where $x_{i}^{(n-1)}=(x_{i}^{(n-2)})^{\prime}$. Hence, starting from degree $5$, symmetric polynomials of $\mathrm{Com}\langle X\rangle^{(D)}/$$\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ are
$x_{1}^{(n-1)}x_{2}x_{3}x_{4}\cdots x_{n}+x_{2}^{(n-1)}x_{1}x_{3}x_{4}\cdots x_{n}+\\ x_{3}^{(n-1)}x_{1}x_{2}x_{4}\cdots x_{n}+\ldots+x_{n}^{(n-1)}x_{1}x_{2}x_{3}\cdots x_{n-1},$
which correspond to $p_{n}$. ∎
4. Free metabelian Lie-admissible algebra
An algebra is called metabelian Lie-admissible if it satisfies the following identities:
$(ab)c-(ba)c-c(ab)+c(ba)+(bc)a-(cb)a-a(bc)\\ +a(cb)+(ca)b-(ac)b-b(ca)+b(ac)=0,$ (5)
$(ab)(cd)=0.$ (6)
Let us consider the polarization of metabelian Lie-admissible algebra, i.e., we consider an algebra with two operations which is defined on metabelian Lie-admissible algebra as follows:
$[a.b]=ab-ba,\;\{a,b\}=ab+ba.$
In this case, the defining identities of the variety of metabelian Lie-admissible algebras become to
$[a,b]=-[b,a],\;\{a,b\}=\{b,a\},$
$[[a,b],c]+[[b,c],a]+[[c,a],b]=0,$
$[[a,b],[c,d]]=[[a,b],\{c,d\}]=[\{a,b\},\{c,d\}]=\{\{a,b\},\{c,d\}\}=\\ \{\{a,b\},[c,d]\}=\{[a,b],[c,d]\}=0.$ (7)
As a consequence, we obtain
$[[[a,b],c],d]=[[[a,b],d],c]$ (8)
and
[[{$[[\{a,b\},c],d]=[[\{a,b\},d],c]$ (9)
which hold in free metabelian Lie-admissible algebra.
Let us construct a basis of the free metabelian Lie-admissible algebra in terms of binary trees with two types of vertices $\bullet$ and $\circ$. We consider only trees of the following form:
$A=$**
$\ldots$
*$x_{i_{1}}$$x_{i_{2}}$$x_{i_{n-1}}$$x_{i_{n}}$
We place on vertices of the tree $\bullet$ and $\circ$ in all possible ways. On the leaves of the tree, we place generators from the countable set $X$. This tree in a unique way corresponds to the sequence $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}}).$
4.1 Example.
If
[…]
then $A$ corresponds to $(\bullet_{x_{1}},\circ_{x_{2}},x_{3})$. We define a set of sequences as follows:
1) If there are several consecutive black dots in a sequence, then all corresponding generators of these vertices are ordered, i.e., for $(\ldots,\circ_{i_{k-1}},\bullet_{i_{k}},\bullet_{i_{k+1}},\ldots,\bullet_{i_{l-1}},\circ_{i_{l}},\ldots)$, we have $i_{k}\geq i_{k+1}\geq\ldots\geq i_{l-1}$;
2) If the rightmost vertex is white then the rightmost generator is less than the previous one, i.e., for $(\ldots,\circ_{i_{n-1}},x_{i_{n}})$, we have $i_{n-1}\leq i_{n}$;
3) If a given consecutive sequence of black dots continues to the rightmost vertex and the number of black dots is bigger than $2$, then all the generators of these vertices are ordered and the rightmost generator is bigger than the previous one, i.e., for $(\ldots,\bullet_{i_{k}},\bullet_{i_{k+1}},\ldots,\bullet_{i_{n-1}},x_{i_{n}})$, we have $i_{k}\geq i_{k+1}\geq\ldots\geq i_{n-1}<i_{n}$;
4) In condition $3$ if the number of black dots is not bigger than $2$, then the generators are ordered as in Lyndon-Shirshov words, i.e., the basis monomials of the free Lie algebra of degrees 2 and 3;
[…]
4.2 Theorem.
The set of monomials $\mathcal{M}$ is a basis of the free metabelian Lie-admissible algebra.
Proof.
[…]
Now, let us consider a free algebra $A\langle X\rangle$ with the basis $\mathcal{M}$. The multiplications $\bullet$ and $\circ$ in $A\langle X\rangle$ are defined as follows:
$\begin{gathered}\begin{cases}X_{1}*X_{2}=0\;\;\text{if $X_{1},X_{2}\in\mathcal{M}$, $\mathrm{deg}(X_{1}),\mathrm{deg}(X_{2})>1$ and $*$ is $\bullet$ or $\circ$};\\ (x_{i_{1}}*(\cdots*(x_{i_{n-1}}*x_{i_{n}})\cdots))\circ x_{j}=x_{j}\circ(x_{i_{1}}*(\cdots*(x_{i_{n-1}}*x_{i_{n}})\cdots)),\\ (x_{i_{1}}\circ(\cdots*(x_{i_{n-1}}*x_{i_{n}})\cdots))\bullet x_{j}=-x_{j}\bullet(x_{i_{1}}\circ(\cdots*(x_{i_{n-1}}*x_{i_{n}})\cdots)),\\ \text{where $n\geq 3$.}\\ (x_{i_{1}}\bullet(\cdots\bullet(x_{i_{k}}\bullet(x_{i_{k+1}}\circ(\cdots*(x_{i_{n-1}}*x_{i_{n}})\cdots)))\cdots))\bullet x_{j}=-x_{i_{1}}\bullet(\cdots\bullet(x_{l}\\ \bullet(x_{j}\bullet(x_{l+1}\bullet(\cdots\bullet(x_{i_{k}}\bullet(x_{i_{k+1}}\circ(\cdots*(x_{i_{n-1}}*x_{i_{n}})\cdots)))\cdots))))\cdots),\\ \text{where $l\leq j\leq l+1$ and $n\geq 3$.}\\ \end{cases}\end{gathered}$
If the monomials $X_{1}$ and $X_{2}$ do not involve $\circ$ then we rewrite the product $X_{1}\bullet X_{2}$ according to the multiplication table of free metabelian Lie algebras.
[…]
Calculating the dimension of operad $\mathrm{M}\mathrm{Lie}$ by means of the package [7], we get the following result:
$n$ 1 2 3 4 5 6 7
$\dim(\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}(n))$ 1 2 11 77 679 7184 88668
We see that the dimension of this operad is growing at a high rate, and this sequence does not coincide with any sequence from OEIS. In [23] was given the dimension of the Lie-admissible operad up to degree $7$, which is
$n$ 1 2 3 4 5 6 7
$\dim(\mathrm{Lie}\textrm{-}\mathrm{adm}(n))$ 1 2 11 101 1299 21484 434314
It will be interesting to find a general formula for the dimension of the metabelian Lie-admissible and Lie-admissible operads.
5. Symmetric polynomials of the free metabelian Lie-admissible algebra
For $n$, let us define the sequence $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})$, where * can be $\circ$ or $\bullet$. For each sequence, we define a space $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$ as follows:
$(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}=\sum_{i_{1},\ldots,i_{n}}x_{i_{1}}*(x_{i_{2}}*(\ldots(x_{i_{n-1}}*x_{i_{n}})\ldots)).$
5.1 Example.
$(\bullet_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\circ)}=\sum_{i_{1},i_{2},i_{3}}x_{i_{1}}\bullet(x_{i_{2}}\circ x_{i_{3}}).$
For each space $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$, $n\geq 3$, we define a polynomial $p_{(*,*,\ldots,*,n)}$ as follows:
1) We divide this sequence into consecutive vertices of the same colour;
2) For $k_{i}$ consecutive white vertices, we select $k_{i}$ generators and write all possible permutations for them;
3) For $k_{i}$ consecutive black vertices, we select $k_{i}$ generating ones and write them in descending order;
4) If the sequence ends with $k_{i}$ vertices of white colour, then for the selected generator we write all possible permutations of $k_{i}$ generators so that the last two generators are always ordered;
5) If the sequence ends with $k_{i}$ vertices of black colour, then for the selected generator we write symmetric polynomials of the free metabelian Lie algebra on $k_{i+1}$ generators.
The sum of such monomials gives polynomial $p_{(*,*,\ldots,*,n)}$. For $n=1,2$, we set $p_{(1)}=x_{1}+x_{2}+\ldots+x_{n}$ and $p_{(2)}=\sum_{i\neq j}x_{i}\circ x_{j}$.
5.2 Example.
For multilinear part of $\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}\langle X\rangle$, let us construct $(\bullet_{x_{i_{1}}},\circ_{x_{i_{2}}},$ $x_{i_{3}}{)}_{(\bullet,\circ)}$, ${(\circ_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})}_{(\bullet,\circ)}$ and ${(\circ_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})}_{(\bullet,\circ)}$.
$p_{(\bullet,\circ,3)}=x_{1}\bullet(x_{2}\circ x_{3})+x_{2}\bullet(x_{1}\circ x_{3})+x_{3}\bullet(x_{1}\circ x_{2}),$
$p_{(\circ,\bullet,3)}=x_{1}\circ(x_{2}\bullet x_{3})+x_{2}\circ(x_{1}\bullet x_{3})+x_{3}\circ(x_{1}\bullet x_{2}),$
$p_{(\circ,\circ,3)}=x_{1}\circ(x_{2}\circ x_{3})+x_{2}\circ(x_{1}\circ x_{3})+x_{3}\circ(x_{1}\circ x_{2}).$
For ${(\bullet_{x_{i_{1}}},\bullet_{x_{i_{2}}},\circ_{x_{i_{3}}},x_{i_{4}})}_{(\bullet,\bullet,\circ)}$, we have
$p_{(\bullet,\bullet,\circ,3)}=x_{2}\bullet(x_{1}\bullet(x_{3}\circ x_{4}))+x_{3}\bullet(x_{1}\bullet(x_{2}\circ x_{4}))+x_{4}\bullet(x_{1}\bullet(x_{2}\circ x_{3}))+\\ x_{3}\bullet(x_{2}\bullet(x_{1}\circ x_{4}))+x_{4}\bullet(x_{2}\bullet(x_{1}\circ x_{3}))+x_{4}\bullet(x_{3}\bullet(x_{1}\circ x_{2})).$
[…]
5.3 Theorem.
For the multilinear part of the free metabelian Lie-admissible algebra, the symmetric polynomials have the form $p_{(*,*,\ldots,*,n)}$.
Proof.
For $n=1,2$, the result is obvious. From the multiplication table of the free metabelian Lie-admissible algebra, one obtains
$\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}_{\geq 3}\langle X\rangle=\oplus_{(*,*,\ldots,*)}(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$
as a vector space, where $\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}_{\geq 3}\langle X\rangle$ is a multilinear part of the free metabelian Lie-admissible algebra of degree greater than $2$. For example, if $n=3$ then
$\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}_{3}\langle X\rangle=(\bullet_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\bullet)}\oplus(\bullet_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\circ)}\oplus\\ (\circ_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\circ,\bullet)}\oplus(\circ_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\circ,\circ)}.$
Moreover, each monomial of the space $(*_{x_{i_{1}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$ under action $S_{n}$ belongs to the same space. It remains to note that $p_{(*,*,\ldots,*,n)}$ is a symmetric polynomial, and for each space $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},$ $x_{i_{n}})_{(*,*,\ldots,*)}$ there is one unique symmetric polynomial which is $p_{(*,*,\ldots,*,n)}$. ∎
ACKNOWLEDGMENTS
This work was supported by UAEU grant G00004614.